# Reward-Safety Balance in Offline Safe RL via Diffusion Regularization

### v1

@@ Abstract
Constrained reinforcement learning (RL) seeks high-performance policies under safety constraints. We focus on an offline setting where the agent has only a fixed dataset—common in realistic tasks to prevent unsafe exploration. To address this, we propose Diffusion-Regularized Constrained Offline Reinforcement Learning (DRCORL), which first uses a diffusion model to capture the behavioral policy from offline data and then extracts a simplified policy to enable efficient inference. We further apply gradient manipulation for safety adaptation, balancing the reward objective and constraint satisfaction. This approach leverages high-quality offline data while incorporating safety requirements. Empirical results show that DRCORL achieves reliable safety performance, fast inference, and strong reward outcomes across robot learning tasks. Compared to existing safe offline RL methods, it consistently meets cost limits and performs well with the same hyperparameters, indicating practical applicability in real-world scenarios.
@@ 1 Introduction
Offline reinforcement learning (RL) has advanced decision-making by learning from pre-collected datasets without online interaction (Fujimoto et al., 2019; Levine et al., 2020). For real-world control tasks (e.g., autonomous driving, industrial control), safety is equally critical. Safe RL addresses this by imposing constraints, often formulated as a constrained Markov decision process (CMDP) (Gu et al., 2024c; Altman, 2021), to ensure high performance without violating safety requirements. These can be hard constraints (no violation at each step) (Zheng et al., 2024; Ganai et al., 2024) or soft constraints (expected total cost below a threshold) (Chow et al., 2018; Yang et al., 2020a). We focus on the soft-constraint setting in this work.
@@ 2 Preliminary
A CMDP (Altman, 2021) is defined by the tuple $\langle\mathcal{S},\mathcal{A},P,r,c,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state and action spaces, respectively. The transition kernel $P$ specifies the probability ${P}(s^{\prime}|s,a)$ of transitioning from state $s$ to state $s^{\prime}$ when action $a$ is taken. The reward function is $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and the cost function is $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The discount factor is denoted by $\gamma$. A policy is a function $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that represents agent’s decision rule, i.e., the agent takes action $a$ with probability $\pi(a|s)$ in state $s$, and we define $\Pi$ as the set of all feasible policies. Under policy $\pi$, the value function is defined as $V^{\pi}_{\diamond}(\rho)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\diamond(s_{t},a_{t})|s_{0}\sim\rho]$, where $\diamond\in\{r,c\}$, $\rho$ is the initial distribution, and the expectation is taken over all possible trajectories. Similarly, the associated Q-function is defined as $Q^{\pi}_{\diamond}(s,a)=\diamond(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}_{\diamond}(s^{\prime})]$ for $\diamond\in\{r,c\}$. In standard CMDP, the objective is to find a policy $\pi\in\Pi$ that maximizes the cumulative rewards $V^{\pi}_{r}(\rho)$ while ensuring that the cumulative cost $V^{\pi}_{c}(\rho)$ remains within a predefined budget $l$.
In the offline setting, the agent cannot interact directly with the environment and instead relies solely on a static dataset consisting of multiple transition tuples $\mathcal{D}^{\mu}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime},c_{i})\}_{i=1}^{N}$, which is collected using a behavioral policy $\pi_{b}(a|s)$. This offline nature introduces the risk of distributional shift between the dataset and the learned policy. To address this, an additional constraint is often imposed to limit the deviation of the learned policy $\pi$ from the behavioral policy $\pi_{b}$, resulting in the optimization problem
@@
where $D_{\operatorname{KL}}(p\|q)$ the KL-divergence between two distributions defined as $D_{\operatorname{KL}}(p\|q)=\mathbb{E}_{x\sim p}\left[\log(p(x)/q(x))\right]$. We use the KL-divergence to penalize the learned policy $\pi$’s distance to the behavioral policy, though it is actually not a distance measure. To address safety constraints, primal-dual-based methods Ding et al. (2021); Paternain et al. (2022); Wu et al. (2024) typically reformulate the constrained optimization problem as follows:
@@ 3.1 Diffusion Model for Policy Extraction
where $w(t)$ is the weight function. Then the reversing denoising process can be formulated as $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\big{(}x_{t}-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\psi}(x_{t},t)\big{)}+\sqrt{\beta_{t}}z_{t}$, where $z_{t}\sim\mathcal{N}(0,I)$. Using this notion, we can similarly use the diffusion model to diffuse in the action space $\mathcal{A}$ and sample actions given the current state with the reversing process. In Wang et al. (2022), the authors learned diffusion policies from the offline dataset using guided diffusion. The diffusion policy here is $\pi_{\psi}(a|s)=\mathcal{N}(a_{T};0,I)\prod_{t=1}^{T}p_{\psi}(a_{t-1}|a_{t},s)$, where $p_{\psi}(a_{t-1}|a_{t},s)$ is a Gaussian distribution with mean $m_{\psi}(a_{t},t|s)$ and variance $\Sigma_{\psi}(a_{t},t|s)$ See also Lu et al. (2023); Hansen-Estruch et al. (2023), where diffusion policy is used for inference in policy evaluation. The shortage of these methods is that the diffusion models are slow in inference speed, even under the improved sampling scheme in Song et al. (2020a); Lu et al. (2022) the reverse denoising process takes at least 10 steps. Therefore, in this work we mainly use diffusion models for pretraining and learning the behavioral policy $\pi_{b}$. For state-action pair $(s,a)$ in the offline dataset $\mathcal{D}^{\mu}$, we train our diffusion policy model by minimizing the loss
@@ 3.2 Diffusion Regularization
The work Chen et al. (2023) showed under the offline RL setting that one can train a simple policy using the pretrained critic and diffusion policy imitating the behavioral policy. The key step is to use the reverse KL divergence to regularize the target policy to be close to the behavioral policy. The forward KL penalizes the KL-divergence $D_{\operatorname{KL}}(\pi^{*}(\cdot|s)\|\pi_{\theta}(\cdot|s))$ while the reverse KL penalizes $D_{\operatorname{KL}}(\pi_{\theta}(\cdot|s)\|\pi^{*}(\cdot|s))$. As shown in Chen et al. (2023), the forward KL leads to modes covering issue while backward KL encourages mode-seeking behavior, although the latter is harder to optimize. Therefore, in this work we also choose the reverse KL for regularization. We also constrain our policy family to a simple Gaussian policy class $\Pi:\{\pi_{\theta}(a|s)=\mathcal{N}(a;m_{\theta}(s),\Sigma_{\theta}(s))\}$. Now, we denote the learned diffusion model’s score function as $\epsilon_{\psi}(a_{t},t|s)$ and the corresponding diffusion policy as $\mu_{\psi}(a|s)$. Then, the reverse KL between the policy $\pi_{\theta}(a|s)$ and the approximated behavioral policy $\mu_{\psi}(a|s)$ can be written as $-\mathcal{H}(\pi_{\theta}(\cdot|s))+H(\pi_{\theta}(\cdot|s),\mu_{\psi}(\cdot|s))$. For Gaussian policy, the first part self-entropy term $\mathcal{H}$ can be directly computed in closed form. For $\mathcal{A}=\mathbb{R}^{d}$, we have
$\mathcal{H}(\pi_{\theta}(\cdot|s))$ $=\int_{\mathcal{A}}-\log\pi_{\theta}(a|s)\pi_{\theta}(a|s)da$ (6)
$=\frac{1}{2}\log(\operatorname{det}(\Sigma_{\theta}(s)))+\frac{d}{2}\log(2\pi)+\frac{d}{2}.$
@@
$H(\pi_{\theta}(\cdot|s),\mu_{\psi}(\cdot|s))$ $=\mathbb{E}_{\pi_{\theta}(\cdot|s)}[-\log\mu_{\psi}(a|s)]$ (7)
$=\mathbb{E}_{z\sim\mathcal{N}(0,I)}[-\log\mu_{\psi}(m_{\theta}+\Sigma_{\theta}^{1/2}z|s)].$
@@
$\nabla_{\theta}D_{\operatorname{KL}}(\pi_{\theta}(\cdot|s)||\mu_{\psi}(\cdot|s))$
$=$ $\nabla_{\theta}H(\pi_{\theta}(\cdot|s),\mu_{\psi}(\cdot|s))-\nabla_{\theta}\mathcal{H}(\pi_{\theta}(\cdot|s))$ (8)
$=$ $\mathbb{E}_{z\sim\mathcal{N}(0,I)}\big{[}-\nabla_{a}\log\mu_{\psi}(m_{\theta}(s)+\Sigma_{\theta}^{1/2}z|s)\cdot\nabla_{\theta}(m_{\theta}(s)+\Sigma_{\theta}^{1/2}z)\big{]}-\frac{1}{2}\nabla_{\theta}\log(\operatorname{det}\Sigma_{\theta}(s)).$
@@ 3.3 Safe Adaptation
To design our algorithm, we define the reward and cost optimization obejectives respectively as follows:
[…]
We consider the reverse KL divergence term in Eq. (10) and Eq. (11) as a regularization term to penalize the policy for deviating too far from the behavioral policy $\pi_{b}$. In practical implementation, we use the learned diffusion policy $\mu_{\psi}$ to replace $\pi_{b}$ and compute the gradient for optimization using Eq. (8). As discussed in Gu et al. (2024a), to address the conflict between optimizing reward and cost, it is essential to directly handle the gradient at each optimization step. We also adopt the gradient manipulation idea in Gu et al. (2024a), but we do so under the offline setting where we no longer have access to updating our critic with the on-policy data. The gradient manipulation method can be generally described as follows. For each step where we need to optimize the reward, we update our gradient with $\theta\leftarrow\theta+\eta g_{r}$, and when we need to optimize the cost our gradient is updated with $\theta\leftarrow\theta+\eta g_{c}$. The gradient manipulation aims at updating the parameter $\theta$ with a linear combination of the two gradients as $\beta_{r}g_{r}+\beta_{c}g_{c}$. We can use the angle $\phi:=\cos^{-1}\left(\frac{\langle g_{r},g_{c}\rangle}{\|g_{r}\|\|g_{c}\|}\right)$ between the two gradients to judge whether the gradients are conflicting or not, namely if $\phi>90^{\circ}$ the two gradients are conflicting and otherwise they are not conflicting. Especially, the ultimate gradient $g$ is computed via the rule
@@
where $g_{r}^{+}=g_{r}-\frac{g_{r}\cdot g_{c}}{\|g_{c}\|^{2}}g_{c}$ is $g_{c}$’s projection on the null space of $g_{r}$ and $g_{c}^{+}=g_{c}-\frac{g_{c}\cdot g_{r}}{\|g_{r}\|^{2}}g_{r}$ is $g_{r}$’s projection on the null space of $g_{c}$. It is also shown in Gu et al. (2024a) that under the assumption of the convexity of the optimization target, one can ensure monotonic improvement using gradient manipulation. As a result, here we also employ the gradient manipulation method to update our policy parameter.
With the procedures outlined above, we present our main algorithm in Algorithm 1, where the safe adaptation step is visualized in Figure 1 and detailed in Algorithm 3 given in Appendix B. Furthermore, we provide the following theoretical results derived from our algorithm, where the proof is deferred to Appendix A.1.
Algorithm 1 DROCORL
1:Input: Dataset $\mathcal{D}^{\mu}$, slack bounds $h^{+}$, $h^{-}$
2:Pretrain diffusion model $\epsilon_{\psi}$
3:Pretrain $\pi_{\theta}$ with diffusion model $\epsilon_{\psi}$
4://Behavior Cloning
5:Pretrain $Q_{r}^{\pi}$, $Q_{c}^{\pi}$
6://Pretrain critic
7:for each gradient step $t$ do
8:     Sample mini-batch =$\mathcal{B}=\{(s_{t},a_{t},s_{t}^{\prime},r_{t},c_{t})\}_{t=1}^{|\mathcal{B}|}$
9:     Update critics for reward and cost
10:     Update diffusion model $\epsilon_{\psi}$ with $\mathcal{B}$
11:     $g=\operatorname{SafeAdaptation}(\pi_{\theta},\epsilon_{\psi},Q^{\pi_{\theta}}_{r},Q^{\pi_{\theta}}_{c},h^{+},h^{-})$
12:     Update $\theta\leftarrow\theta+\eta g$
13:end for
@@
Under the tabular MDP setting, the policy follows a softmax parameterization, where the policy $\pi$ is parameterized by $\theta:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The policy is expressed as $\pi(a|s)=\frac{\exp(\theta(s,a))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\theta(s,a^{\prime}))}$.
Policy improvement is performed using the natural policy gradient method (Kakade and Langford, 2002), where the policy parameter $\theta$ is updated at each iteration as
$\theta\leftarrow\theta+\eta\big{(}\mathcal{F}^{\theta}_{\rho}\big{)}^{\dagger}\nabla_{\theta}V^{\pi}(\rho),$ (14)
where $\mathcal{F}^{\theta}_{\rho}=\mathbb{E}_{s\sim\rho,a\sim\pi}\big{[}\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}\big{]}$ denotes the Fisher information matrix, and the $\dagger$ operator denotes the Moore-Penrose inverse. Based on this setup, we now present the convergence guarantees.
@@
$\mathbb{E}[V^{\tilde{\pi}}_{c}(\rho)]-b\leq\mathcal{O}(\epsilon_{\text{offline}})+\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}T}}\right),$ (16)

### v2

# Don’t Trade Off Safety: Diffusion Regularization for Constrained Offline RL

@@ Abstract
Constrained reinforcement learning (RL) seeks high-performance policies under safety constraints. We focus on an offline setting where the agent learns from a fixed dataset—a common requirement in realistic tasks to prevent unsafe exploration. To address this, we propose Diffusion-Regularized Constrained Offline Reinforcement Learning (DRCORL), which first uses a diffusion model to capture the behavioral policy from offline data and then extracts a simplified policy to enable efficient inference. We further apply gradient manipulation for safety adaptation, balancing the reward objective and constraint satisfaction. This approach leverages high-quality offline data while incorporating safety requirements. Empirical results show that DRCORL achieves reliable safety performance, fast inference, and strong reward outcomes across robot learning tasks. Compared to existing safe offline RL methods, it consistently meets cost limits and performs well with the same hyperparameters, indicating practical applicability in real-world scenarios.
@@ 1 Introduction
Offline reinforcement learning (RL) has advanced decision-making by learning from pre-collected datasets without online interaction (Fujimoto et al., 2019; Levine et al., 2020). For real-world control tasks (e.g., autonomous driving, industrial control), safety is equally critical. Safe RL addresses this by imposing constraints, often formulated as a constrained Markov decision process (CMDP) (Gu et al., 2024c; Altman, 2021), to ensure high performance without violating safety requirements. These can be hard constraints (no violation at each step) (Zheng et al., 2024; Ganai et al., 2024) or soft constraints (expected total cost below a threshold) (Chow et al., 2018; Yang et al., 2020; Koirala et al., 2024; Guo et al., 2025; Shi et al., 2021). We focus on the soft-constraint setting in this work.
@@ 2 Preliminary
A Constrained Markov Decision Process (CMDP) (Altman, 2021) is defined by the tuple $\langle\mathcal{S},\mathcal{A},P,r,c,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state and action spaces, respectively. The transition kernel $P$ specifies the probability ${P}(s^{\prime}|s,a)$ of transitioning from state $s$ to state $s^{\prime}$ when action $a$ is taken. The reward function is $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and the cost function is $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The discount factor is denoted by $\gamma$. A policy is a function $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that represents the agent’s decision rule, i.e., the agent takes action $a$ with probability $\pi(a|s)$ in state $s$, and we define $\Pi$ as the set of all feasible policies. Under policy $\pi$, the value function is defined as $V^{\pi}_{\diamond}(\rho)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\diamond(s_{t},a_{t})|s_{0}\sim\rho]$, where $\diamond\in\{r,c\}$, $\rho$ is the initial distribution, and the expectation is taken over all possible trajectories. Similarly, the associated Q-function is defined as $Q^{\pi}_{\diamond}(s,a)=\diamond(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}_{\diamond}(s^{\prime})]$ for $\diamond\in\{r,c\}$. In standard CMDP, the objective is to find a policy $\pi\in\Pi$ that maximizes the cumulative rewards $V^{\pi}_{r}(\rho)$ while ensuring that the cumulative cost $V^{\pi}_{c}(\rho)$ remains within a predefined budget $l$.
In the offline setting, the agent cannot interact directly with the environment and instead relies solely on a static dataset $\mathcal{D}^{\mu}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime},c_{i})\}_{i=1}^{N}$ consisting of multiple transition tuples, which is collected using a behavioral policy $\pi_{b}(a|s)$. This offline nature introduces the risk of distributional shift between the dataset and the learned policy. To mitigate this, an additional constraint is often imposed to limit the deviation of the learned policy $\pi$ from the behavioral policy $\pi_{b}$, resulting in the optimization problem
@@
where $D_{\operatorname{KL}}(p\|q)$ the KL-divergence between two distributions defined as $D_{\operatorname{KL}}(p\|q)=\mathbb{E}_{x\sim p}\left[\log(p(x)/q(x))\right]$. We use the KL-divergence to penalize the learned policy $\pi$’s distance to the behavioral policy, though it is actually not a distance measure. To address safety constraints, primal-dual methods Ding et al. (2021); Paternain et al. (2022); Wu et al. (2024) typically reformulate the constrained optimization problem as follows:
@@ 3.1 Diffusion Model for Policy Extraction
where $w(t)$ is the weight function. Then the reversing denoising process can be formulated as $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\big{(}x_{t}-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\psi}(x_{t},t)\big{)}+\sqrt{\beta_{t}}z_{t}$, where $z_{t}\sim\mathcal{N}(0,I)$. Using this notion, we can similarly use the diffusion model to diffuse in the action space $\mathcal{A}$ and sample actions given the current state with the reversing process. In Wang et al. (2022), the authors learned diffusion policies from the offline dataset using guided diffusion. The diffusion policy here is $\pi_{\psi}(a|s)=\mathcal{N}(a_{T};0,I)\prod_{t=1}^{T}p_{\psi}(a_{t-1}|a_{t},s)$, where $p_{\psi}(a_{t-1}|a_{t},s)$ is a Gaussian distribution with mean $m_{\psi}(a_{t},t|s)$ and variance $\Sigma_{\psi}(a_{t},t|s)$ . See also Lu et al. (2023); Hansen-Estruch et al. (2023), where diffusion policy is used for inference in policy evaluation. The limitation of these methods is that the diffusion models are slow in inference speed, even under the improved sampling scheme in Song et al. (2020a); Lu et al. (2022) the reverse denoising process takes at least 10 steps. Therefore, in this work we mainly use diffusion models for pretraining and learning the behavioral policy $\pi_{b}$. For state-action pair $(s,a)$ in the offline dataset $\mathcal{D}^{\mu}$, we train our diffusion policy model by minimizing the loss
@@ 3.2 Diffusion Regularization
The work Chen et al. (2023) showed under the offline RL setting that one can train a simple policy using the pretrained critic and diffusion policy imitating the behavioral policy. The key step is to use the reverse KL divergence to regularize the target policy to be close to the behavioral policy. The forward KL is the KL-divergence $D_{\operatorname{KL}}(\pi^{*}(\cdot|s)\|\pi_{\theta}(\cdot|s))$ while the reverse KL is $D_{\operatorname{KL}}(\pi_{\theta}(\cdot|s)\|\pi^{*}(\cdot|s))$. As shown in Chen et al. (2023), the forward KL leads to mode covering issue while backward KL encourages mode-seeking behavior, although the latter is harder to optimize. Therefore, in this work we also choose the reverse KL for regularization. We also constrain our policy family to a simple Gaussian policy class $\Pi:\{\pi_{\theta}(a|s)=\mathcal{N}(a;m_{\theta}(s),\Sigma_{\theta}(s))\}$. Now, we denote the learned diffusion model’s score function as $\epsilon_{\psi}(a_{t},t|s)$ and the corresponding diffusion policy as $\mu_{\psi}(a|s)$. Then, the reverse KL between the policy $\pi_{\theta}(a|s)$ and the approximated behavioral policy $\mu_{\psi}(a|s)$ can be written as $\mathcal{H}(\pi_{\theta}(\cdot|s))+H(\pi_{\theta}(\cdot|s),\mu_{\psi}(\cdot|s))$. For Gaussian policy, the first part self-entropy term $\mathcal{H}$ can be directly computed in closed form. For $\mathcal{A}=\mathbb{R}^{d}$, we have
$\mathcal{H}(\pi_{\theta}(\cdot|s))$ $=\int_{\mathcal{A}}-\log\pi_{\theta}(a|s)\pi_{\theta}(a|s)da=\frac{1}{2}\log(\operatorname{det}(\Sigma_{\theta}(s)))+\frac{d}{2}\log(2\pi)+\frac{d}{2}.$ (6)
@@
$H(\pi_{\theta}(\cdot|s),\mu_{\psi}(\cdot|s))$ $=\mathbb{E}_{\pi_{\theta}(\cdot|s)}[-\log\mu_{\psi}(a|s)]=\mathbb{E}_{z\sim\mathcal{N}(0,I)}[-\log\mu_{\psi}(m_{\theta}+\Sigma_{\theta}^{1/2}z|s)].$ (7)
@@
$\nabla_{\theta}D_{\operatorname{KL}}(\pi_{\theta}(\cdot|s)||\mu_{\psi}(\cdot|s))=\nabla_{\theta}H(\pi_{\theta}(\cdot|s),\mu_{\psi}(\cdot|s))-\nabla_{\theta}\mathcal{H}(\pi_{\theta}(\cdot|s))$ (8)
[…]
@@ 3.3 Safe Adaptation
To design our algorithm, we define the reward and cost optimization objectives respectively as follows:
$\text{Reward:}\max_{\pi\in\Pi}\mathbb{E}_{s\sim\mathcal{D}^{\mu}}\big{[}$ $V^{\pi}_{r}(s)-\frac{1}{\beta}D_{\operatorname{KL}}(\pi(\cdot|s)||\pi_{b}(\cdot|s))\big{]}$ (10)
$\text{Cost:}\max_{\pi\in\Pi}\mathbb{E}_{s\sim\mathcal{D}^{\mu}}\big{[}$ $-(V^{\pi}_{c}(s)-l)-\frac{1}{\beta}D_{\operatorname{KL}}(\pi(\cdot|s)||\pi_{b}(\cdot|s))\big{]}.$ (11)
We consider the reverse KL divergence term in Eq. (10) and Eq. (11) as a regularization term to penalize the policy for deviating too far from the behavioral policy $\pi_{b}$. In practice implementation, we use the learned diffusion policy $\mu_{\psi}$ to replace $\pi_{b}$ and compute the gradient for optimization using Eq. (8). As discussed in Gu et al. (2024a), to address the conflict between optimizing reward and cost, it is essential to directly handle the gradient at each optimization step. We also adopt the gradient manipulation idea in Gu et al. (2024a), but we do so under the offline setting where we no longer have access to updating our critic with the on-policy data. The gradient manipulation method can be generally described as follows. For each step where we need to optimize the reward, we update our gradient with $\theta\leftarrow\theta+\eta g_{r}$, and when we need to optimize the cost our gradient is updated with $\theta\leftarrow\theta+\eta g_{c}$. The gradient manipulation aims at updating the parameter $\theta$ with a linear combination of the two gradients as $\beta_{r}g_{r}+\beta_{c}g_{c}$. We can use the angle $\phi:=\cos^{-1}\left(\frac{\langle g_{r},g_{c}\rangle}{\|g_{r}\|\|g_{c}\|}\right)$ between the two gradients to judge whether the gradients are conflicting or not, namely if $\phi>90^{\circ}$ the two gradients are conflicting and otherwise they are not conflicting. Especially, the final gradient $g$ is computed via the rule
@@
where $g_{r}^{+}=g_{r}-\frac{g_{r}\cdot g_{c}}{\|g_{c}\|^{2}}g_{c}$ is $g_{c}$’s projection on the null space of $g_{r}$ and $g_{c}^{+}=g_{c}-\frac{g_{c}\cdot g_{r}}{\|g_{r}\|^{2}}g_{r}$ is $g_{r}$’s projection on the null space of $g_{c}$. In Gu et al. (2024a) under the assumption of the convexity of the optimization target, one can ensure monotonic improvement using gradient manipulation. Therefore we also employ the gradient manipulation method to update our policy. With the procedures outlined above, we present our Algorithm 1, where the safe adaptation step is visualized in Figure 1 and detailed in Algorithm 2 given in Appendix B. Furthermore, we provide the following theoretical results derived from our algorithm, where the proof is deferred to Appendix A.1.
@@
Algorithm 1 DRCORL
1:Input: Dataset $\mathcal{D}^{\mu}$, slack bounds $^{+}, hh^{-}$
2:Pretrain diffusion model $\epsilon_{\psi}$
3:// Behavior Cloning
4:Pretrain $Q_{r}^{\pi}$, $Q_{c}^{\pi}$
5:// Pretrain critic
6:for each gradient step $t$ do
7:  Sample mini-batch $\mathcal{B}$
8:  Update critics for reward and cost
9:  $g\leftarrow\operatorname{SafeAdaptation}(\dots)$
10:  Update $\theta\leftarrow\theta+\eta g$
11:end for
@@
Under the tabular MDP setting, the policy follows a softmax parameterization, where the policy $\pi$ is parameterized by $\theta:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The policy is expressed as $\pi_{\theta}(a|s)=\frac{\exp(\theta(s,a))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\theta(s,a^{\prime}))}$.
We use natural policy gradient method (Kakade and Langford, 2002) to update policy, where the policy parameter $\theta$ is updated as $\theta\leftarrow\theta+\eta\big{(}\mathcal{F}^{\theta}_{\rho}\big{)}^{\dagger}\nabla_{\theta}V^{\pi}(\rho)$ at each iteration, where $\mathcal{F}^{\theta}_{\rho}=\mathbb{E}_{s\sim\rho,a\sim\pi}\big{[}\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}\big{]}$ is the Fisher information matrix, and the $\dagger$ operator is the Moore-Penrose inverse.
@@
$\mathbb{E}[V^{\tilde{\pi}}_{c}(\rho)]-b$ $\leq\mathcal{O}(\epsilon_{\text{offline}})+\mathcal{O}\left(\sqrt{\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{3}T}}\right).$ (15)